OpenStudy (bananas1234):

If f(x) = 9 – 4|x + 1|, find f(–5). (Points : 1) –15 –7 7 23

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle f(x) = 9 - 4\left|x + 1\right| }\) To find f(-5), plug in -5 for x.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle f(\color{red}{-5}) = 9 - 4\left|\color{red}{-5} + 1\right| }\) this is how you plug it in. Now, you only need to simplify the right side, but remember that absolute value (of any number) is always positive.

OpenStudy (bananas1234):

1 for -5 + 1

OpenStudy (bananas1234):

then we would have f(-5) = 9 - 4 /1/

OpenStudy (bananas1234):

what would i do next?

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle f(-5) = 9 - 4|-5+1|}\) \(\large\color{slate}{\displaystyle f(-5) = 9 - 4|-4|}\) \(\large\color{slate}{\displaystyle f(-5) = 9 - 4(4)}\)

OpenStudy (solomonzelman):

see how it is?

OpenStudy (bananas1234):

so then it would be 4 times 4 right? That would equal 16. so we would have 9-16 left?

OpenStudy (bananas1234):

and that would equal -7. So would the answure be B, -7?

OpenStudy (solomonzelman):

yes, -7 is correct.

OpenStudy (bananas1234):

Thank you!